OpenStudy (anonymous):

Calculate the surface area of a square pyramid where the height is 4 feet and the length of a side is 6 feet?

OpenStudy (anonymous):

@mathstudent55 @atreyu6s3x6

OpenStudy (anonymous):

yea? :D

OpenStudy (anonymous):

A=96ft^2 (not sure )

OpenStudy (anonymous):

can yall help?

OpenStudy (anonymous):

okay do you know the equation? @caitlyn10

OpenStudy (anonymous):

is it 96ft2?

OpenStudy (anonymous):

sorry if i reply slow im helping a couple other people too :)

OpenStudy (anonymous):

no your fine

OpenStudy (anonymous):

okay so do you know the equation for surface area?

OpenStudy (anonymous):

thats it up there

OpenStudy (anonymous):

@atreyu6s3x6 the equation is up there

OpenStudy (anonymous):

okay so what are our numbers? a=? h=?

OpenStudy (anonymous):

or is it 84?

OpenStudy (anonymous):

h=4 a=6

OpenStudy (anonymous):

@atreyu6s3x6

OpenStudy (anonymous):

okay so if we plug that in we have: SA=6^2+2(6)\[\sqrt{6^2/4+4^2}\]

OpenStudy (anonymous):

sorry thats all one equation

OpenStudy (anonymous):

53?@atreyu6s3x6

OpenStudy (anonymous):

i don't think i did that right

OpenStudy (anonymous):

so we simplify: \[36+12\sqrt{36/4+16}\]

OpenStudy (anonymous):

try and simplify that... :)

OpenStudy (anonymous):

96?

OpenStudy (anonymous):

YYYYAAAASSSS!!!!! haha good job :D

OpenStudy (anonymous):

okay would it be 96 or 96^2? @atreyu6s3x6

OpenStudy (anonymous):

it would be 96 ft^2

OpenStudy (anonymous):

awesome! thanks

OpenStudy (anonymous):

no problem :)